### v2

# Computing algebraic Belyi functions on Bring’s curve

Abstract.
In this paper, we explicitly compute two kinds of algebraic Belyi functions on Bring’s curve. One is related to a congruence subgroup of ${\rm SL}_{2}(\mathbb{Z})$ and the other is relate to a congruence subgroup of the triangle group $\Delta(2,4,5)\subset{\rm SL}_{2}(\mathbb{R})$. To carry out the computation, we use elliptic cusp forms of weight 2 for the former case while the automorphism group of Bring’s curve for the latter case. We also discuss a suitable base field (a number field) for isomorphisms between Hulek-Craig’s curve, Bring’s curve, and its another algebraic model obtained as a modular curve.
Key words and phrases:
Belyi functions, Bring’s curve, Hulek-Craig’s curve, modular curves, triangular Shimura curves
1. Introduction
The main result of Belyi [1] shows that a complex projective smooth curve $C$ has a model, as an algebraic curve, defined over a number field if and only if there exists a non-constant function $\beta:C\longrightarrow\mathbb{P}^{1}$ ramified at most over $\{0,1,\infty\}$. In this paper, we call such a function $\beta$ an algebraic Belyi function and more precisely, an algebraic Belyi function defined over a number field $K$ if it is defined over $K$. Thus, any projective smooth geometrically connected curve over a number field is endowed with an algebraic Belyi function. However, finding such a $\beta$ explicitly for a given $C$ is another matter. In fact, many people have studied explicit computation of algebraic Belyi functions (cf. [23],[21] among others). Belyi’s result can also be understood by hyperbolic uniformization and there are two types of such uniformization:
  1. (1)
    (Non-compact case) There exist a finite index subgroup $\Gamma$ of ${\rm SL}_{2}(\mathbb{Z})$ and a biholomorphism $X(\Gamma):=\Gamma\backslash(\mathbb{H}\cup\mathbb{P}^{1}(\mathbb{Q}))\simeq C(\mathbb{C})$ such that $\beta$ is interpreted as the natural projection
    $X(\Gamma)\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{P}^{1}(\mathbb{C}),\ \Gamma\tau\mapsto{\rm SL}_{2}(\mathbb{Z})\tau\mapsto j(\tau)$
    where $j$ is Klein’s $j$-invariant function so that $j(\tau)=q^{-1}+744+196884q+\cdots,\ q:=e^{2\pi\sqrt{-1}\tau},\ \tau\in\mathbb{H}:=\{\tau\in\mathbb{C}\ |\ {\rm Im}(\tau)>0\}$.
  2. (2)
    (Compact case) There exist a finite index subgroup $\Delta$ of some trignale group $\Delta(a,b,c)\subset{\rm SL}_{2}(\mathbb{R})$ and a biholomorphism $X_{\Delta}:=\Delta\backslash\mathbb{H}\simeq C(\mathbb{C})$ such that $\beta$ is interpreted as the natural projection
    $X_{\Delta}\longrightarrow X_{\Delta(a,b,c)}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{P}^{1}(\mathbb{C}),\ \Delta\tau\mapsto\Delta(a,b,c)\tau\mapsto j_{\Delta(a,b,c)}(\tau)$
    where $j_{\Delta(a,b,c)}$ is a uniformizer of $X_{\Delta(a,b,c)}$.
We refer [17, p.71, Theorem] for the non-compact case and [24, Section 3.3, Theorem 3] for the compact case.
In this paper, using the above two interpretation, we will compute two algebraic Belyi functions on Bring’s curve defined by
(1.1) $B:\left\{\begin{array}[]{l}x_{0}+x_{1}+x_{2}+x_{3}+x_{4}=0\\ x^{2}_{0}+x^{2}_{1}+x^{2}_{2}+x^{2}_{3}+x^{2}_{4}=0\\ x^{3}_{0}+x^{3}_{1}+x^{3}_{2}+x^{3}_{3}+x^{3}_{4}=0\end{array}\right.$
inside $\mathbb{P}^{4}$ with homogeneous coordinates $[x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]$. The readers may consult with Theorem 2.2 and Theorem 3.1 as the main results of this paper.
Let us explain about some key points for the main results. Let $S_{5}$ be the symmetric group on 5 letters. It is known that
  1. (1)
    (non-compact case [9, p.82-83]) $B(\mathbb{C})\simeq X(\Gamma_{B})$ where $\Gamma_{B}\subset{\rm SL}_{2}(\mathbb{Z})$ is explicitly given in the next section (see also Section 4), and
  2. (2)
    (compact case [3, p.73]) $\ΔB(\mathbb{C})\simeq\Delta\backslash\mathbb{H}$ for some normal congruence subgroup $\Delta$ of $\Delta(3,4,5)$ such that $\Delta(3,4,5)/\Delta\simeq S_{5}$. Thus, ${\rm Aut}(B(\mathbb{C}))\simeq{\rm Aut}(\Delta\backslash\mathbb{H})=\Delta(3,4,5)/\Delta\simeq S_{5}$.
For the non-compact case, we compute the $j$-function in terms of the modular functions on $X(\Gamma_{B})$ by regarding $j$ and modular functions with algebraic coordinates.
For the compact case, we first use the uniformization to conclude the natural quotient map $B(\mathbb{C})\mapsto B(\mathbb{C})/{\rm Aut}_{\mathbb{C}}(B(\mathbb{C}))$ is an algebraic Belyi function. As a projective smooth complex curve, Bring’s curve is a unique curve of genus 4 with $S_{5}$ as the automorphism group. On the other hands, $S_{5}$ acts properly on the coordinates of $B$ and if we view $B$ as an algebraic curve over $\mathbb{Q}$, we have ${\rm Aut}_{\mathbb{Q}}(B)\supset S_{5}$ where ${\rm Aut}_{\mathbb{Q}}$ stands for the group of all automorphims defined over $\mathbb{Q}$. However, $S_{5}\subset{\rm Aut}_{\mathbb{Q}}(B)\subset{\rm Aut}_{\mathbb{C}}(B(\mathbb{C}))\simeq S_{5}$. Thus, ${\rm Aut}_{\mathbb{Q}}(B)=S_{5}$. Hence, the natural quotient map decent to the quotient map $B\longrightarrow B/{\rm Aut}_{\mathbb{Q}}(B)\simeq\mathbb{P}^{1}_{\mathbb{Q}}$ which is an algebraic Belyi function defined over $\mathbb{Q}$.
We will organize this paper as follows. In Section 2, we study another model of $B$ which is explicitly given as an algebraic model of the modular curve $X(\Gamma_{B})$. By using elliptic cusp forms of weight 2 and theory of canonical curves, we compute an explicit defining equation over $\mathbb{Q}$ for $X(\Gamma_{B})$. In Section 3, we will compute the quotient map $B\mapsto B/{\rm Aut}_{\mathbb{Q}}(B)$ explicitly. Finally, in Section 4, we relate our algebraic model of $X(\Gamma_{B})$ with Hulek-Craig’s curve and discuss a number field $K$ such that these curves are isomorphic over $K$ to Bring’s curve. This shows that Bring’s curve inherits two interesting algebraic Belyi functions.
Acknowledgments. We would like to thank Professors K. Hulek and H. Braden for helpful discussions.
2. Bring’s curve as a modular curve
In this section, we compute the Belyi function on the Bring’s curve $B$ as the modular curve $X(\Gamma_{B})$ introduced as below. It will be explained in Section 4 that $B$ is isomorphic over $\mathbb{Q}(\zeta_{5})$ to an algebraic model over $\mathbb{Q}$ of $X(\Gamma_{B})$. Thus, we may work on $X(\Gamma_{B})$ to compute the Belyi function.
We refer [5] for the basic facts of elliptic modular forms and modular curves. For each positive integer $N$, we denote by $\Gamma(N)\supset\Gamma_{0}(N)\supset\Gamma_{1}(N)$ the three kinds of congruence subgroups inside ${\rm SL}_{2}(\mathbb{Z})$ introduced in [5, p.13, Definition 1.2.1]. For such a congruence subgroup $\Gamma$, we define the (compact) modular curve $X(\Gamma):=\Gamma\backslash(\mathbb{H}\cup\mathbb{P}^{1}(\mathbb{Q}))$ and put $X_{0}(N)=X(\Gamma_{0}(N))$ for simplicity. We denote by $S_{2}(\Gamma)$ the space of all cusp forms of weight 2 with respect to $\Gamma$. It is known that we have an isomorphism $S_{2}(\Gamma)\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{0}(X(\Gamma),\Omega^{1}_{X(\Gamma)}),\ f(\tau)\mapsto f(\tau)d\tau$ where the right hand side stands for the space of all holomorphic 1-forms on $X(\Gamma)$.
2.1. Defining equations over $\mathbb{Q}$ of $X(\Gamma_{B})$
Let
[…]
As explained in [2, Section 2], $B(\mathbb{C})$ is isomorphic to $X(\Gamma_{B})$ as a compact Riemann surface. Notice that $\left(\begin{array}[]{cc}5^{-1}&0\\ 0&1\end{array}\right)\Gamma_{B}\left(\begin{array}[]{cc}5&0\\ 0&1\end{array}\right)=\Gamma_{0}(2,5):=\Gamma_{0}(2)\cap\Gamma(5)$ in the notation there and also $\Gamma_{1}(50)\subset\Gamma_{B}\subset\Gamma_{0}(50)$. Then, the natural inclusion $\Gamma_{B}\subset{\rm SL}_{2}(\mathbb{Z})$ yields the quotient map
(2.1) $\beta_{\Gamma_{B}}:X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{P}^{1}(\mathbb{C}),\ \Gamma_{B}\tau\mapsto{\rm SL}_{2}(\mathbb{Z})\tau\mapsto\frac{1}{1728}j(\tau).$
It is well-known (cf. [6, Proposition 2.1]) that $\beta_{\Gamma_{B}}$ is a Belyi function which is ramified only at $0,1,\infty$. Since each of $X(\Gamma_{B})$ and $X({\rm SL}_{2}(\mathbb{Z}))$ has a structure as a projective smooth algebraic curve defined over $\mathbb{Q}$, the Belyi function $\beta_{\Gamma_{B}}$ can be seen as an algebraic function $\beta^{{\rm alg}}_{\Gamma_{B}}$ over $\mathbb{Q}$. Henceforth, we will explicitly compute $\beta^{{\rm alg}}_{\Gamma_{B}}$.
Let us first compute defining equations over $\mathbb{Q}$ of $X(\Gamma_{B})$.
Proposition 2.1.
The modular curve $X(\Gamma_{B})$ has the following algebraic model over $\mathbb{Q}$ as a projective smooth geometrically connected algebraic curve defined over $\mathbb{Q}:$
(2.2) $X:\left\{\begin{array}[]{l}-(x^{2}+y^{2})+2(z^{2}+w^{2})=0\\ x^{3}+xy^{2}+2y^{3}-4xz^{2}-8xzw=0\end{array}\right.$
inside $\mathbb{P}^{3}$ with homogeneous coordinates $[x:y:z:w]$.
Proof.
[…]
$S_{2}(\Gamma_{B})=\bigoplus_{\chi\in\widehat{(\mathbb{Z}/50\mathbb{Z})^{\times}}\atop\chi^{4}=1}S_{2}(\Gamma_{0}(50),\chi)=S_{2}(\Gamma_{0}(50))\oplus S_{2}(\Gamma_{0}(50),\chi_{2})$
[…]
(2.3) $\begin{array}[]{l}f_{1}(\tau):=f_{{\rm 50.2.a.a}}(q)=q-q^{2}+q^{3}+q^{4}-q^{6}+2*q^{7}-\cdots\\ f_{2}(\tau):=f_{{\rm 50.2.a.b}}(q)=q+q^{2}-q^{3}+q^{4}-q^{6}-2*q^{7}+q^{8}-\cdots\\ f_{3}(\tau):=\frac{1}{2}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})+f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q-q^{4}-q^{6}+\cdots\\ f_{4}(\tau):=\frac{1}{2\sqrt{-1}}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})-f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q^{2}+q^{3}-2q^{7}-q^{8}-\cdots\end{array}$
where each label $\ast$ of $f_{\ast}$ corresponds to the corresponding datum on [13] and we can read off the Fourier coefficients up to $q^{99}$ for which it is enough for the computation below. Since Bring’s curve is non-hyperelliptic and of genus 4, its canonical divisor is very ample. Hence, by noting $S_{2}(\Gamma_{B})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H^{0}(X(\Gamma_{B}),\Omega^{1}_{X(\Gamma_{B})}),\ f(\tau)\mapsto f(\tau)d\tau$, we have the canonical embedding:
$\Phi:X(\Gamma_{B})\hookrightarrow\mathbb{P}^{3}(\mathbb{C}),\ \tau\mapsto[x:y:z:w],\ (x,y,z,w):=(f_{1}(\tau),f_{2}(\tau),f_{3}(\tau),f_{4}(\tau)).$
By using Petri’s theorem as exlpained in [20, Section 2.2] for $X_{0}(N)$ (which also applies to $\Gamma_{B}$), the image of $\Phi$ is given by the intersection of a quadratic equation and a cubic equation in $\mathbb{P}^{3}$. The claim follows by determining the coefficients of those equations by using (2.3) (see [20] for details in the computation). ∎
2.2. An algebraic Belyi function on $X$
In this subsection, we compute the algebraic function over $\mathbb{Q}$ corresponding to the quotient map $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle{\rm SL}_{2}(\mathbb{Z})\tau\mapsto j(\tau)}}{{\simeq}}\mathbb{P}^{1}(\mathbb{C})$ by using the intermediate modular curve $X_{0}(50)$ so that the quotient map factors through it:
$X(\Gamma_{B})\stackrel{{\scriptstyle\pi_{1}}}{{\longrightarrow}}X_{0}(50)\stackrel{{\scriptstyle\pi_{2}}}{{\longrightarrow}}X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle j\atop\sim}}{{\longrightarrow}}\mathbb{P}^{1}(\mathbb{C}).$
Let $X_{0}(50)^{{\rm alg}}$ be the hyperelliptic curve over $\mathbb{Q}$ defined by
(2.4) $t^{2}=s^{6}-4s^{5}-10s^{3}-4s+1$
such that $X_{0}(50)^{{\rm alg}}(\mathbb{C})\simeq X_{0}(50)$ as a compact Riemann surface (see [8, p.290]). An algebraic function over $\mathbb{Q}$ for $j\circ\pi_{2}$ has been computed in [8, p.290] with (2.4) is explicitly given by
$X_{0}(50)^{{\rm alg}}_{\mathbb{Q}}\longrightarrow\mathbb{P}^{1}_{\mathbb{Q}},\ (s,t)\mapsto j=h(s,t):=\frac{A(s)+B(s)t}{2s^{25}(s^{4}-s^{3}+s^{2}-s+1)^{2}}$
where the polynomials $A(x)$ and $B(x)$ in the one variable $x$ are given in [8, p.290]. Thus, we have only to compute an algebraic map between $X$ and (2.4) corresponding to $\pi_{1}$.
Put $T:=y/z=f_{1}/f_{2}$ and $S:=z/w=f_{3}/f_{4}$ so that $T,S\in\mathbb{C}(X_{0}(50))$. By using Gröebner basis for the ideal generated by $zT-y,wS-z$ and (2.2) in $\mathbb{Q}[x,y,z,w,T,S]$, we can eliminate $x,y,z,w$ to obtain a relation $01+S^{2}+2S^{3}-4T-4S^{2}T-T^{2}-S^{2}T^{2}+2S^{3}T^{2}=$ 0. The curve defined by this equation is isomorphic over $\mathbb{Q}$ to (2.4) by
$(S,T)=\Bigg{(}\frac{s+1}{s-1},\frac{s^{3}-s^{2}+s-1+t}{s(s^{2}+s+2)}\Bigg{)},\ (s,t)=\Bigg{(}\frac{S+1}{S-1},\frac{2(-2-2S^{2}+T+S^{2}T+2S^{3}T)}{(S-1)^{3}}\Bigg{)}.$
Thus, we have proved the following theorem:
Theorem 2.2.
The algebraic function
$\beta^{{\rm alg}}_{\Gamma_{B}}:X\longrightarrow\mathbb{P}^{1}_{\mathbb{Q}},\ [x:y:z:w]\mapsto h\Bigg{(}\frac{z+w}{z-w},\frac{2(w^{3}x-2w^{3}y+wxz^{2}-2wyz^{2}+2xz^{3})}{y(z-w)^{3}}\Bigg{)}$
is an algebraic Belyi function defined over $\mathbb{Q}$.
Remark 2.3.
The Belyi function $X(G_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\simeq\mathbb{P}^{1}(\mathbb{C})$ in question is a non-Galois covering of $\mathbb{P}^{1}(\mathbb{C})$. Thus, it gives rise to a non-regular dessin on $X(G_{B})$.
2.3. A decomposition of the Jacobian variety over a number field
In this section, we study a number field $L$ such that the Jacobian variety ${\rm Jac}(X)$ of $X$ decomposes completely over $L$. For two abelian varieties $A_{1},A_{2}$ over a number field $K$ and a field extension $M/K$, we write $A_{1}\stackrel{{\scriptstyle M}}{{\sim}}A_{2}$ if $A_{1}$ is isogenous over $M$ to $A_{2}$. For each newform $f$ in $S_{2}(\Gamma_{B})$, we denote by $A_{f}$ the Shimura’s abelian variety over $\mathbb{Q}$ of dimension $[\mathbb{Q}_{f}:\mathbb{Q}]$ where $\mathbb{Q}_{f}$ is the Hecke field of $f$ (cf. [16, Section 3]). We note that $A_{f}$ is $\mathbb{Q}$-simple.
Proposition 2.4.
Keep the notation being in Proposition 2.2. Let $L=\mathbb{Q}(\sqrt{5},\sqrt{5+2\sqrt{5}})$. Then, it holds that
${\rm Jac}(X)\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}A_{f_{{\rm 50.2.b.a}}}\times J_{0}(50)_{\mathbb{Q}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}A_{f_{{\rm 50.2.b.a}}}\times E_{{\rm 50.a3}}\times E_{{\rm 50.b3}}\stackrel{{\scriptstyle L}}{{\sim}}E^{4}$
where $E_{\ast}$ is the elliptic curve over $\mathbb{Q}$ with conductor 50 whose label $\ast$ corresponds to one in [13] and $E=E_{{\rm 50.a3}}:y^{2}+xy+y=x^{3}-x-2$ is an elliptic curve over $\mathbb{Q}$ with $j$-invariant $-\frac{25}{2}$.
Proof.
The decomposition over $\mathbb{Q}$ follows from [15, Proposition (2.3)]. It follows from [13] that $A_{f_{{\rm 50.2.b.a}}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}{\rm Res}_{\mathbb{Q}(\sqrt{5})/\mathbb{Q}}(E_{1})$ where $E_{1}=E_{100.1-{\rm a}3}:y^{2}+(\varepsilon+1)xy+(\varepsilon+1)y=x^{3}+\varepsilon x^{2},\ \varepsilon:=\displaystyle\frac{1+\sqrt{5}}{2}$ with the $j$-invariant $-\frac{25}{2}$. Here ${\rm Res}_{\mathbb{Q}(\sqrt{5})/\mathbb{Q}}$ stands for the Weil-restriction from $\mathbb{Q}(\sqrt{5})$ to $\mathbb{Q}$. Since $E_{1}$ and $E$ are isomorphic each other over $\overline{\mathbb{Q}}$, by [22, Chapter III, Proposition 3.1], there exist $u,r,s,t\in\overline{\mathbb{Q}}$ such that $E\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}E_{1},\ (x,y)\mapsto(u^{2}x+r,u^{3}y+u^{2}sx+r)$. Solving the system of the algebraic equations, we have
$s^{2}+s+\varepsilon^{2}=0,\ r=-\varepsilon,\ u=\varepsilon^{-2}(1+2s).$
On the other hand, by [22, Chapter III, Proposition 3.1] again, $E=E_{{\rm 50.a3}}\stackrel{{\scriptstyle\mathbb{Q}(\sqrt{5})}}{{\simeq}}E_{{\rm 50.b3}},\ (x,y)\mapsto(u^{\prime 2}x+r^{\prime},u^{\prime 3}y+u^{\prime 2}s^{\prime}x+r^{\prime})$ with $r^{\prime}=0,\ s^{\prime}=\displaystyle\frac{1}{10}(-5+\sqrt{5}),\ t^{\prime}=\frac{1}{50}(-25+\sqrt{5}),\ u^{\prime}=\frac{\sqrt{5}}{5}$. Thus, the claim follows. ∎
Remark 2.5.
Since $A_{f_{{\rm 50.2.b.a}}}$ is a $\mathbb{Q}$-simple abelian surface, ${\rm Jac}(X)$ never be isogenous over $\mathbb{Q}$ to the product of four elliptic curves over $\mathbb{Q}$.
3. Bring’s curve as a triangular (Shimura) curve
Let $\Delta(2,4,5)\subset{\rm SL}_{2}(\mathbb{R})$ be the triangle group corresponding to the triple $(a,b,c)=(2,4,5)$ (see [3, Section 2]) so that $F=\mathbb{Q}(\sqrt{2},\sqrt{5})$ and $E=\mathbb{Q}(\sqrt{5})$ in the notation there. By [3, Proposition 5.13], there exists an embedding $\Delta(2,4,5)\hookrightarrow{\rm SL}_{2}(\mathcal{O}_{E})$ where $\mathcal{O}_{E}$ stands for the ring of inters of $E$. We define $\Delta$ by the kernel of $\Delta(2,4,5)\longrightarrow{\rm SL}_{2}(\mathcal{O}_{E})\longrightarrow{\rm SL}_{2}(\mathcal{O}_{E}/(\sqrt{5}))$ and it yields the isomorphism ${\rm Aut}(X_{\Delta})=\Delta(2,4,5)/\Delta\simeq{\rm SL}_{2}(\mathcal{O}_{E}/(\sqrt{5}))\simeq{\rm SL}_{2}(\mathbb{F}_{5})\simeq S_{5}$. To be more precise, $\Delta(2,4,5)/\{\pm 1\}$ can be regarded as the group of the reduced norm 1 of a maximal order inside the quaternionic algebra $A:=\Big{(}\frac{-1,\varepsilon}{E}\Big{)}$ over $E$ where $\varepsilon=\frac{1+\sqrt{5}}{2}$ (see [3, p.53]). Thus, $X_{\Delta(2,4,5)}=(\Delta(2,4,5)/\{\pm 1\})\backslash\mathbb{H}$ is a Shimura curve (of genus zero). Therefore, we have a $S_{5}$-Galois covering $X_{\Delta}\longrightarrow X_{\Delta(2,4,5)}$ between triangular Shimura curves which is a Belyi function by construction. The quotient map $B\longrightarrow B/{\rm Aut}_{\mathbb{Q}}(B)=B/S_{5}\simeq\mathbb{P}^{1}_{\mathbb{Q}}$ is an algebraic Belyi function over $\mathbb{Q}$ since its base change to $\mathbb{C}$ is corresponding to the quotient map $X_{\Delta(2,4,5)}\longrightarrow X_{\Delta}$. Thus, we have only to find out a generator of the function field of $B/S_{5}$. Let $\rho:B\longrightarrow B$ and $\kappa:B\longrightarrow B$ be the automorphism corresponding to $(12345),(12)\in S_{5}$ respectively. Notice that $\rho,\kappa$ generate $S_{5}$. Put $T=(x_{1}x_{2}x_{3}x_{4})/x^{4}_{5}$ and
(3.1) $t:=\sum_{i=1}^{5}\rho^{i}(T)+\sum_{i=1}^{5}\kappa\rho^{i}(T).$
By direct computation, we can check $t$ is invariant under the action of $S_{5}$.
Theorem 3.1.
Keep the notation being as above. The rational function $t$ is a generator of $\mathbb{Q}(B/S_{5})$. It gives rise to an algebraic Belyi function over $\mathbb{Q}$:
$B\longrightarrow\mathbb{P}^{1},\ [x_{1}:x_{2},x_{3}:x_{4}:x_{5}]\mapsto t.$
Proof.
Put $v:=x_{1}/x_{0}$ for simplicity. By using Grobner basis, we have $t=\displaystyle\frac{-2g(v)}{v^{4}(1+v)^{4}(1+v^{2})^{4}}$ where
$g(v)=1+5v+15v^{2}+35v^{3}+65v^{4}+101v^{5}+135v^{6}+155v^{7}+165v^{8}+165v^{9}+161v^{10}$
$+165v^{11}+165v^{12}+155v^{13}+135v^{14}+101v^{15}+65v^{16}+35v^{17}+15v^{18}+5v^{19}+v^{20}.$
For a generic $t\in\mathbb{Q}$, we have the number of $v$ satisfying the above equation is 20. Further, for each $v$, we have $6$ pairs of $(x_{2}/x_{5},x_{3}/x_{5})$ modulo the equality $\sum_{i=1}^{5}x_{i}=0$. Thus, $[\mathbb{Q}(B):\mathbb{Q}(t)]=20\times 6=120$. On the other hand, we observe that $[\mathbb{Q}(B):\mathbb{Q}(B/S_{5})]=[\mathbb{C}(B):\mathbb{C}(B/S_{5})]=|\Delta(2,3,5)/\Delta|=|S_{5}|=120$ and $\mathbb{Q}(t)\subset\mathbb{Q}(B/S_{5})\subset\mathbb{Q}(B)$. Thus, the claim follows. ∎
Remark 3.2.
Since any (compact) Shimura curve does not posses any real points (cf. [19]), $t$ can not generate the function field of any rational model over $E=\mathbb{Q}(\sqrt{5})$ of $X_{\Delta(2,4,5)}$.
4. A relation between two algebraic models
In this section, we study the field of definition of isomorphisms between $B$, $X$ and the Hulek-Craig curve. It is known that Bring’s curve $B$ is isomorphic to the algebraic model $X$ of $X(\Gamma_{B})$ over $\overline{\mathbb{Q}}$. We would like to discuss a number field $K$ such that $B$ is isomorphic over $K$ to $X$. Obviously, $K$ cannot be $\mathbb{Q}$ since $B(\mathbb{Q})=\emptyset$ while $X(\mathbb{Q})\neq\emptyset$. For another reason, it is known that ${\rm Jac}(B)\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}E^{4}$ (cf. [18, p.87, Exercise-b)]) where $E$ is given in Proposition 2.4. Thus, the field $K$ cannot be $\mathbb{Q}$ since ${\rm Jac}(B)$ is not isogenous over $\mathbb{Q}$ to ${\rm Jac}(X)$ (see also Remark 2.5). We should note that ${\rm Jac}(X)$ is isogenous over $L$ to ${\rm Jac}(B)$, but it is not necessarily true that $X$ is isomorphic over $L$ to $B$. Thus, we need more careful study for our purpose.
However, fortunately, it is known from [2, Proposition 2.9 and its proof] that $B$ is isomorphic over $\mathbb{Q}(\zeta_{5})$ to Hulek-Craig’s curve $X^{{\rm HC}}$ which is given by the desingularization of a projective geometrically irreducible singular curve
$F(x_{1},y_{1},z_{1}):=x_{1}(y^{5}_{1}+z^{5}_{1})+x^{2}_{1}y^{2}_{1}z^{2}_{1}-x^{4}_{1}y_{1}z_{1}-2y_{1}^{3}z^{3}_{1}=0$
inside $\mathbb{P}^{2}$ with the coordinates $[x_{1}:y_{1}:z_{1}]$. This curve has studied in [9] and [4]. As explained in [4, p.27] that Hulek-Craig’s curve has the parametrization in terms of modular forms:
$x_{1}=\sum_{n\in\mathbb{Z}}(-q)^{(5n)^{2}},\ y_{1}=\sum_{n\in\mathbb{Z}}(-q)^{(5n+1)^{2}},\ z_{1}=\sum_{n\in\mathbb{Z}}(-q)^{(5n+2)^{2}},\ q=e^{2\pi\sqrt{-1}\tau}$
where we slightly modified $q$ with $-q$ from Craig’s notation. In terms of the Jacobi’s theta function $\theta_{a,b}(z,\tau)$, $(z,\tau)\in\mathbb{C}\times\mathbb{H}$ with the rational character $(a,b)\in\mathbb{Q}^{2}$ (see [14, p.10]), we see that
$x_{1}=\theta_{0,0}(0,25(2\tau+1)),\ y_{1}=\theta_{\frac{1}{5},0}(0,25(2\tau+1)),\ z_{1}=\theta_{\frac{2}{5},0}(0,25(2\tau+1))$
and thus, $x_{1},y_{1},z_{1}$ are modular forms of weight $\frac{1}{2}$ with respect to the double covering of $\Gamma(5)\cap\Gamma_{1}(4)$ (see [12, Chapter IV]). It is known (cf. [2, Lemma 2.6]) that
$v_{1}=\frac{(Y^{3}-X)dX}{F_{Y}(X,Y,1)},\ v_{2}=\frac{(Y^{2}X-1)dX}{F_{Y}(X,Y,1)},\ v_{3}=\frac{(Y-X^{2})dX}{F_{Y}(X,Y,1)},\ v_{4}=\frac{Y(X^{2}-Y)dX}{F_{Y}(X,Y,1)}$
where $X=x_{1}/z_{1},\ Y=y_{1}/z_{1}$, and $F_{Y}(X,Y,1):=\frac{\partial F(X,Y,1)}{\partial Y}$. Then, in terms of $q$-expansion, we have
$(v_{1},v_{2},v_{3},v_{4})=(-q^{3}+q^{8}+4q^{13}\cdots,q^{2}-2q^{7}-q^{12}\cdots,-q^{4}+2q^{9}+2q^{14}\cdots,-q+q^{6}+3q^{11}\cdots)dq$
$=(f_{1},f_{2},f_{3},f_{4})A(dq),\ A=\left(\begin{array}[]{cccc}-\frac{1}{4}&-\frac{1}{4}&-\frac{1}{4}&-\frac{1}{4}\\ \frac{1}{4}&\frac{1}{4}&-\frac{1}{4}&-\frac{1}{4}\\ 0&0&\frac{1}{2}&-\frac{1}{2}\\ -\frac{1}{2}&\frac{1}{2}&0&0\end{array}\right)$
Put $w_{i}:=F_{Y}(X,Y,1)\displaystyle\frac{v_{i}}{dX}$ for $i=1,2,3,4$ and let us regard $[f_{1}dq:f_{2}dq:f_{3}dq:f_{4}dq]$ with the coordinates $[x:y:z:w]$ of $X$ in view of the theory of canonical curves. Then, $[x:y:z:w]=[w_{1}:w_{2}:w_{3}:w_{4}]A^{-1}$. Using this, we have a birational map $\psi$ from the singular model of $X^{{\rm HC}}$ defined by $F(x,y,z)$ to $X$ given by $\psi([x_{1}:y_{1}:z_{1}])=$
(4.1) $\psi([x_{1}:y_{1}:z_{1}])=[x:y:z:w]$
where
$\left\{\begin{array}[]{l}x=(z_{1}-y_{1})(x^{2}_{1}+x_{1}y_{1}+y^{2}_{1}+x_{1}z_{1}+z^{2}_{1}),\\ y=(y_{1}-z_{1})(-x^{2}_{1}+x_{1}y_{1}+y^{2}_{1}+x_{1}z_{1}+2y_{1}z_{1}+z^{2}_{1}),\\ z=(y_{1}+z_{1})(-x^{2}_{1}+y_{1}z_{1}),\\ w=x_{1}y^{2}_{1}-y^{3}_{1}+x_{1}z^{2}_{1}-z^{3}_{1}.\end{array}\right.$
Then, it yields an isomorphism over $\mathbb{Q}$ from $X^{{\rm HC}}$ to $X$. This result is not obvious from the construction explained in [9, p.82-83]. Thus, we have checked the following result:
Theorem 4.1.
Recall that the projective smooth model $X$ over $\mathbb{Q}$ of the modular curve $X(\Gamma_{B})$ from (2.2). It holds that
  1. (1)
    Bring’s curve $B$ is isomorphic over $\mathbb{Q}(\zeta_{5})$ to $X^{{\rm HC}}$ whose isomorphism is given by [2, Proposition 2.9], and
  2. (2)
    Hulek-Craig’s curve $X^{{\rm HC}}$ is isomorphic over $\mathbb{Q}$ to $X$ whose isomorphism is given by (4.1). In particular, Hulek-Craig’s curve is isomorphic over $\mathbb{Q}$ to a canonical model of $X(\Gamma_{B})$ (see [11] for canonical models of modular curves).
Thus, $B$ inherits two algebraic Belyi functions over $\mathbb{Q}(\zeta_{5})$ such that one is non-regular (not a Galois covering) while the other is regular (a Galois covering) (see [7] for the definition of regular Belyi functions).

### v3

@@
 In this paper, we explicitly compute two kinds of algebraic Belyi functions on Bring’s curve.
-One is related to a congruence subgroup of ${\rm SL}_{2}(\mathbb{Z})$ and the other is relate to a congruence subgroup of the triangle group $\Delta(2,4,5)\subset{\rm SL}_{2}(\mathbb{R})$.
-To carry out the computation, we use elliptic cusp forms of weight 2 for the former case while the automorphism group of Bring’s curve for the latter case.
-We also discuss a suitable base field (a number field) for isomorphisms between Hulek-Craig’s curve, Bring’s curve, and its another algebraic model obtained as a modular curve.
+One is related to a congruence subgroup of ${\rm SL}_{2}(\mathbb{Z})$ and the other is related to a congruence subgroup of the triangle group $\Delta(2,4,5)\subset{\rm SL}_{2}(\mathbb{R})$.
+To carry out the computation, we use elliptic cusp forms of weight 2 for the former case and the automorphism group of Bring’s curve for the latter case.
+We also discuss a suitable base field (a number field) for describing isomorphisms between Hulek-Craig’s curve, Bring’s curve, and another algebraic model obtained as a modular curve.
 Key words and phrases:
@@
 However, finding such a $\beta$ explicitly for a given $C$ is another matter.
-In fact, many people have studied explicit computation of algebraic Belyi functions (cf.
+In fact, the explicit computation of algebraic Belyi functions has been studied extensively (cf.
 [23],[21] among others).
-Belyi’s result can also be understood by hyperbolic uniformization and there are two types of such uniformization:
+Belyi’s result can also be understood via hyperbolic uniformization and there are two types of uniformization:
   1. (1)
-    (Non-compact case) There exist a finite index subgroup $\Gamma$ of ${\rm SL}_{2}(\mathbb{Z})$ and a biholomorphism $X(\Gamma):=\Gamma\backslash(\mathbb{H}\cup\mathbb{P}^{1}(\mathbb{Q}))\simeq C(\mathbb{C})$ such that $\beta$ is interpreted as the natural projection
-    $X(\Gamma)\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{P}^{1}(\mathbb{C}),\ \Gamma\tau\mapsto{\rm SL}_{2}(\mathbb{Z})\tau\mapsto j(\tau)$
-    where $j$ is Klein’s $j$-invariant function so that $j(\tau)=q^{-1}+744+196884q+\cdots,\ q:=e^{2\pi\sqrt{-1}\tau},\ \tau\in\mathbb{H}:=\{\tau\in\mathbb{C}\ |\ {\rm Im}(\tau)>0\}$.
+    (Non-compact case) There exists a finite index subgroup $\Gamma$ of ${\rm SL}_{2}(\mathbb{Z})$ and a biholomorphism $X(\Gamma):=\Gamma\backslash(\mathbb{H}\cup\mathbb{P}^{1}(\mathbb{Q}))\simeq C(\mathbb{C})$ such that $\beta$ is interpreted as the natural projection
+    $X(\Gamma)\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{P}^{1}(\mathbb{C}),\ \Gamma\tau\mapsto{\rm SL}_{2}(\mathbb{Z})\tau\mapsto\frac{1}{1728}j(\tau)$
+    where $j$ is Klein’s $j$-invariant function, with expansion $j(\tau)=q^{-1}+744+196884q+\cdots,\ q:=e^{2\pi\sqrt{-1}\tau},\ \tau\in\mathbb{H}:=\{\tau\in\mathbb{C}\ |\ {\rm Im}(\tau)>0\}$.
   2. (2)
-    (Compact case) There exist a finite index subgroup $\Delta$ of some trignale group $\Delta(a,b,c)\subset{\rm SL}_{2}(\mathbb{R})$ and a biholomorphism $X_{\Delta}:=\Delta\backslash\mathbb{H}\simeq C(\mathbb{C})$ such that $\beta$ is interpreted as the natural projection
+    (Compact case) There exists a finite index subgroup $\Delta$ of some triangle group $\Delta(a,b,c)\subset{\rm SL}_{2}(\mathbb{R})$ and a biholomorphism $X_{\Delta}:=\Delta\backslash\mathbb{H}\simeq C(\mathbb{C})$ such that $\beta$ is interpreted as the natural projection
     $X_{\Delta}\longrightarrow X_{\Delta(a,b,c)}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbb{P}^{1}(\mathbb{C}),\ \Delta\tau\mapsto\Delta(a,b,c)\tau\mapsto j_{\Delta(a,b,c)}(\tau)$
     where $j_{\Delta(a,b,c)}$ is a uniformizer of $X_{\Delta(a,b,c)}$.
-We refer [17, p.71, Theorem] for the non-compact case and [24, Section 3.3, Theorem 3] for the compact case.
-In this paper, using the above two interpretation, we will compute two algebraic Belyi functions on Bring’s curve defined by
+We refer to [17, p.71, Theorem] for the non-compact case and [24, Section 3.3, Theorem 3] for the compact case.
+In this paper, using the above two interpretations, we compute two algebraic Belyi functions on Bring’s curve defined by
 (1.1) $B:\left\{\begin{array}[]{l}x_{0}+x_{1}+x_{2}+x_{3}+x_{4}=0\\ x^{2}_{0}+x^{2}_{1}+x^{2}_{2}+x^{2}_{3}+x^{2}_{4}=0\\ x^{3}_{0}+x^{3}_{1}+x^{3}_{2}+x^{3}_{3}+x^{3}_{4}=0\end{array}\right.$
-inside $\mathbb{P}^{4}$ with homogeneous coordinates $[x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]$. The readers may consult with Theorem 2.2 and Theorem 3.1 as the main results of this paper.
-Let us explain about some key points for the main results. Let $S_{5}$ be the symmetric group on 5 letters. It is known that
+inside $\mathbb{P}^{4}$ with homogeneous coordinates $[x_{0}:x_{1}:x_{2}:x_{3}:x_{4}]$.
+The symmetric group $S_{5}$ of degree 5 acts on $B$ by permutation of coordinates and it is known that $S_{5}$ is the full automorphism group of $B$.
+We refer to [25, Example 5.5 and Proposition 5.6] for another aspect of Bring’s curve.
+A key ingredient in computing the Belyi maps is to relate Bring’s curve with a certain modular curve and Hulek-Craig’s curve.
+We briefly introduce these curves and the details are given in the corresponding sections.
+Let $X(\Gamma_{B})^{{\rm alg}}$ be the algebraic model over $\mathbb{Q}$ of the compactified modular curve $X(\Gamma_{B}):=\overline{\Gamma_{B}\backslash\mathbb{H}}$ which is given by
+(1.2) $X(\Gamma_{B})^{{\rm alg}}:\left\{\begin{array}[]{l}-(x^{2}+y^{2})+2(z^{2}+w^{2})=0\\ x^{3}+xy^{2}+2y^{3}-4xz^{2}-8xzw=0\end{array}\right.$
+inside $\mathbb{P}^{4}$ (see Section 2).
+Here $\Gamma_{B}:=\Gamma_{0}(2)\cap\Gamma(5)$ is a congruence subgroup in ${\rm SL}_{2}(\mathbb{Z})$.
+Let $\beta^{{\rm alg}}_{\Gamma_{B}}:X(\Gamma_{B})^{{\rm alg}}\longrightarrow\mathbb{P}^{1}_{\mathbb{Q}}$ be an algebraic model over $\mathbb{Q}$ of $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\simeq\mathbb{P}^{1}(\mathbb{C}),\tau\mapsto\frac{1}{1728}j$ (see Theorem 2.2).
+On the other hand, Hulek-Craig’s curve $X^{{\rm HC}}$ over $\mathbb{Q}$ is defined by the desingularization of a projective geometrically irreducible singular curve
+$F(x_{1},y_{1},z_{1}):=x_{1}(y^{5}_{1}+z^{5}_{1})+x^{2}_{1}y^{2}_{1}z^{2}_{1}-x^{4}_{1}y_{1}z_{1}-2y_{1}^{3}z^{3}_{1}=0$
+inside $\mathbb{P}^{2}$ with the coordinates $[x_{1}:y_{1}:z_{1}]$.
+This curve has studied in [9] and [4].
+Hulek-Craig’s curve $X^{{\rm HC}}$ appears in the context of the Horrocks-Mumford bundle [9].
+Through its symmetries, the latter is also closely related to the group $\Gamma(5)$.
+Then we prove the following result:
+Theorem 1.1.
+(Theorem 4.1) Keep the notation as above. Let $\zeta_{5}$ be a primitive fifth root of unity.
   1. (1)
-    (non-compact case [9, p.82-83]) $B(\mathbb{C})\simeq X(\Gamma_{B})$ where $\Gamma_{B}\subset{\rm SL}_{2}(\mathbb{Z})$ is explicitly given in the next section (see also Section 4), and
+    Bring’s curve $B$ is isomorphic over $\mathbb{Q}(\zeta_{5})$ to Hulek-Craig’s curve $X^{{\rm HC}}$;
+  2. (2)
+    Hulek-Craig’s curve $X^{{\rm HC}}$ is isomorphic over $\mathbb{Q}$ to $X(\Gamma_{B})^{{\rm alg}}$.
+We note that the first part of Theorem 1.1 is due to [2, Proposition 2.9]. The main results of this paper are Theorem 2.2 and Theorem 3.1 and they are summarized as follows:
+Theorem 1.2.
+Bring curve has two algebraic Belyi functions which are given by the following contexts:
+  1. (1)
+    (non-compact type) The composition of $B\stackrel{{\scriptstyle\mathbb{Q}(\zeta_{5})}}{{\simeq}}X^{{\rm HC}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\simeq}}X(\Gamma_{B})^{{\rm alg}}$ and $\beta^{{\rm alg}}_{\Gamma_{B}}:X(\Gamma_{B})^{{\rm alg}}\longrightarrow\mathbb{P}^{1}_{\mathbb{Q}}$ gives an algebraic Belyi map over $\mathbb{Q}(\zeta_{5})$ on $B$.
+  2. (2)
+    (compact type) The natural quotient $B\longrightarrow B/S_{5}$ gives an algebraic Belyi map over $\mathbb{Q}$ on $B$.
+It is an algebraic model over $\mathbb{Q}$ of the natural quotient map $X_{\Delta}\longrightarrow X_{\Delta(3,4,5)}\simeq\mathbb{P}^{1}(\mathbb{C})$ induced by the inclusion of a certain subgroup $\Delta$ of $\Delta(3,4,5)$.
+Let us explain about some key points for the main results. It is known that
+  1. (1)
+    (non-compact case [9, p.82-83]) $B(\mathbb{C})\simeq X(\Gamma_{B})$ as a Riemann surface, and
   2. (2)
@@
 Thus, ${\rm Aut}(B(\mathbb{C}))\simeq{\rm Aut}(\Delta\backslash\mathbb{H})=\Delta(3,4,5)/\Delta\simeq S_{5}$.
-For the non-compact case, we compute the $j$-function in terms of the modular functions on $X(\Gamma_{B})$ by regarding $j$ and modular functions with algebraic coordinates.
-For the compact case, we first use the uniformization to conclude the natural quotient map $B(\mathbb{C})\mapsto B(\mathbb{C})/{\rm Aut}_{\mathbb{C}}(B(\mathbb{C}))$ is an algebraic Belyi function.
-As a projective smooth complex curve, Bring’s curve is a unique curve of genus 4 with $S_{5}$ as the automorphism group.
+For the non-compact case, we compute an algebraic model of the $j$-function in terms of the modular functions on $X(\Gamma_{B})$ by regarding $j$ and modular functions with algebraic coordinates.
+For the compact case, we first use the uniformization to conclude that the natural quotient map $B(\mathbb{C})\mapsto B(\mathbb{C})/{\rm Aut}_{\mathbb{C}}(B(\mathbb{C}))$ is an algebraic Belyi function.
+As a projective smooth complex curve, Bring’s curve is the unique curve of genus 4 with $S_{5}$ as the automorphism group.
 On the other hands, $S_{5}$ acts properly on the coordinates of $B$ and if we view $B$ as an algebraic curve over $\mathbb{Q}$, we have ${\rm Aut}_{\mathbb{Q}}(B)\supset S_{5}$ where ${\rm Aut}_{\mathbb{Q}}$ stands for the group of all automorphims defined over $\mathbb{Q}$.
@@
 In Section 2, we study another model of $B$ which is explicitly given as an algebraic model of the modular curve $X(\Gamma_{B})$.
-By using elliptic cusp forms of weight 2 and theory of canonical curves, we compute an explicit defining equation over $\mathbb{Q}$ for $X(\Gamma_{B})$.
+By using elliptic cusp forms of weight 2 and the theory of canonical curves, we compute an explicit defining equation over $\mathbb{Q}$ for $X(\Gamma_{B})$.
 In Section 3, we will compute the quotient map $B\mapsto B/{\rm Aut}_{\mathbb{Q}}(B)$ explicitly.
@@
 This shows that Bring’s curve inherits two interesting algebraic Belyi functions.
-Acknowledgments. We would like to thank Professors K. Hulek and H. Braden for helpful discussions.
+Acknowledgments.
+We would like to thank Professors K.
+Hulek and H.
+Braden for helpful discussions.
+We also thank Natalia Amburg for kindly informing us an article [25].
+Finally, we would like to give special thanks to the referee, whose suggestions have greatly improved the presentation and readability of this paper.
 2. Bring’s curve as a modular curve
-In this section, we compute the Belyi function on the Bring’s curve $B$ as the modular curve $X(\Gamma_{B})$ introduced as below.
+In this section, we compute the Belyi function on Bring’s curve $B$ as the modular curve $X(\Gamma_{B})$ defined below.
 It will be explained in Section 4 that $B$ is isomorphic over $\mathbb{Q}(\zeta_{5})$ to an algebraic model over $\mathbb{Q}$ of $X(\Gamma_{B})$.
 Thus, we may work on $X(\Gamma_{B})$ to compute the Belyi function.
-We refer [5] for the basic facts of elliptic modular forms and modular curves.
+We refer to [5] for the basic facts of elliptic modular forms and modular curves.
 For each positive integer $N$, we denote by $\Gamma(N)\supset\Gamma_{0}(N)\supset\Gamma_{1}(N)$ the three kinds of congruence subgroups inside ${\rm SL}_{2}(\mathbb{Z})$ introduced in [5, p.13, Definition 1.2.1].
@@
 It is well-known (cf.
-[6, Proposition 2.1]) that $\beta_{\Gamma_{B}}$ is a Belyi function which is ramified only at $0,1,\infty$.
+[6, Proposition 2.1]) that $\beta_{\Gamma_{B}}$ is a Belyi function ramified only at $0,1,\infty$.
 Since each of $X(\Gamma_{B})$ and $X({\rm SL}_{2}(\mathbb{Z}))$ has a structure as a projective smooth algebraic curve defined over $\mathbb{Q}$, the Belyi function $\beta_{\Gamma_{B}}$ can be seen as an algebraic function $\beta^{{\rm alg}}_{\Gamma_{B}}$ over $\mathbb{Q}$.
@@
 The modular curve $X(\Gamma_{B})$ has the following algebraic model over $\mathbb{Q}$ as a projective smooth geometrically connected algebraic curve defined over $\mathbb{Q}:$
-(2.2) $X:\left\{\begin{array}[]{l}-(x^{2}+y^{2})+2(z^{2}+w^{2})=0\\ x^{3}+xy^{2}+2y^{3}-4xz^{2}-8xzw=0\end{array}\right.$
+(2.2) $X(\Gamma_{B})^{{\rm alg}}:\left\{\begin{array}[]{l}-(x^{2}+y^{2})+2(z^{2}+w^{2})=0\\ x^{3}+xy^{2}+2y^{3}-4xz^{2}-8xzw=0\end{array}\right.$
 inside $\mathbb{P}^{3}$ with homogeneous coordinates $[x:y:z:w]$.
@@
 […]
-(2.3) $\begin{array}[]{l}f_{1}(\tau):=f_{{\rm 50.2.a.a}}(q)=q-q^{2}+q^{3}+q^{4}-q^{6}+2*q^{7}-\cdots\\ f_{2}(\tau):=f_{{\rm 50.2.a.b}}(q)=q+q^{2}-q^{3}+q^{4}-q^{6}-2*q^{7}+q^{8}-\cdots\\ f_{3}(\tau):=\frac{1}{2}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})+f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q-q^{4}-q^{6}+\cdots\\ f_{4}(\tau):=\frac{1}{2\sqrt{-1}}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})-f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q^{2}+q^{3}-2q^{7}-q^{8}-\cdots\end{array}$
+(2.3) $\begin{array}[]{l}f_{1}(\tau):=f_{{\rm 50.2.a.a}}(q)=q-q^{2}+q^{3}+q^{4}-q^{6}+2q^{7}-\cdots\\ f_{2}(\tau):=f_{{\rm 50.2.a.b}}(q)=q+q^{2}-q^{3}+q^{4}-q^{6}-2q^{7}+q^{8}-\cdots\\ f_{3}(\tau):=\frac{1}{2}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})+f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q-q^{4}-q^{6}+\cdots\\ f_{4}(\tau):=\frac{1}{2\sqrt{-1}}(f_{{\rm 50.2.b.a}}(q,\sqrt{-1})-f_{{\rm 50.2.b.a}}(q,-\sqrt{-1}))=q^{2}+q^{3}-2q^{7}-q^{8}-\cdots\end{array}$
 where each label $\ast$ of $f_{\ast}$ corresponds to the corresponding datum on [13] and we can read off the Fourier coefficients up to $q^{99}$ for which it is enough for the computation below.
+Readers may visit the website https://www.lmfdb.org/ModularForm/GL2/Q/holomorphic/ linked in [13] and enter 50 for the level, and 2 for the weight, and either 1 or 2 for the character order.
+Then all labelled forms as above will show up individually.
 Since Bring’s curve is non-hyperelliptic and of genus 4, its canonical divisor is very ample.
@@
 $\Phi:X(\Gamma_{B})\hookrightarrow\mathbb{P}^{3}(\mathbb{C}),\ \tau\mapsto[x:y:z:w],\ (x,y,z,w):=(f_{1}(\tau),f_{2}(\tau),f_{3}(\tau),f_{4}(\tau)).$
-By using Petri’s theorem as exlpained in [20, Section 2.2] for $X_{0}(N)$ (which also applies to $\Gamma_{B}$), the image of $\Phi$ is given by the intersection of a quadratic equation and a cubic equation in $\mathbb{P}^{3}$.
-The claim follows by determining the coefficients of those equations by using (2.3) (see [20] for details in the computation).
+By using Petri’s theorem as explained in [20, Section 2.2] for $X_{0}(N)$ (which also applies to $\Gamma_{B}$), the image of $\Phi$ is given by the intersection of a quadratic equation and a cubic equation in $\mathbb{P}^{3}$.
+The claim follows by determining the coefficients of those equations by using (2.3); see [20] for computational details.
 ∎
-2.2. An algebraic Belyi function on $X$
-In this subsection, we compute the algebraic function over $\mathbb{Q}$ corresponding to the quotient map $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle{\rm SL}_{2}(\mathbb{Z})\tau\mapsto j(\tau)}}{{\simeq}}\mathbb{P}^{1}(\mathbb{C})$ by using the intermediate modular curve $X_{0}(50)$ so that the quotient map factors through it:
+2.2. An algebraic Belyi function on $X(\Gamma_{B})^{{\rm alg}}$
+In this subsection, we compute the algebraic function over $\mathbb{Q}$ corresponding to the quotient map $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle{\rm SL}_{2}(\mathbb{Z})\tau\mapsto\frac{1}{1728}j(\tau)}}{{\simeq}}\mathbb{P}^{1}(\mathbb{C})$ by using the intermediate modular curve $X_{0}(50)$ so that the quotient map factors through it:
 $X(\Gamma_{B})\stackrel{{\scriptstyle\pi_{1}}}{{\longrightarrow}}X_{0}(50)\stackrel{{\scriptstyle\pi_{2}}}{{\longrightarrow}}X({\rm SL}_{2}(\mathbb{Z}))\stackrel{{\scriptstyle j\atop\sim}}{{\longrightarrow}}\mathbb{P}^{1}(\mathbb{C}).$
+Here $\pi_{1}$ and $\pi_{2}$ are natural projections induced by the inclusion $\Gamma_{B}\subset\Gamma_{0}(50)\subset{\rm SL}_{2}(\mathbb{Z})$.
 Let $X_{0}(50)^{{\rm alg}}$ be the hyperelliptic curve over $\mathbb{Q}$ defined by
@@
 Put $T:=y/z=f_{1}/f_{2}$ and $S:=z/w=f_{3}/f_{4}$ so that $T,S\in\mathbb{C}(X_{0}(50))$.
-By using Gröebner basis for the ideal generated by $zT-y,wS-z$ and (2.2) in $\mathbb{Q}[x,y,z,w,T,S]$, we can eliminate $x,y,z,w$ to obtain a relation $01+S^{2}+2S^{3}-4T-4S^{2}T-T^{2}-S^{2}T^{2}+2S^{3}T^{2}=$ 0.
+By using Mathematica version 12.1, we can compute a Gröebner basis for the ideal generated by $zT-y,wS-z$ and (2.2) in $\mathbb{Q}[x,y,z,w,T,S]$ such that we can find a relation $01+S^{2}+2S^{3}-4T-4S^{2}T-T^{2}-S^{2}T^{2}+2S^{3}T^{2}=$ 0 in that basis.
 The curve defined by this equation is isomorphic over $\mathbb{Q}$ to (2.4) by
@@
 The algebraic function
-$\beta^{{\rm alg}}_{\Gamma_{B}}:X\longrightarrow\mathbb{P}^{1}_{\mathbb{Q}},\ [x:y:z:w]\mapsto h\Bigg{(}\frac{z+w}{z-w},\frac{2(w^{3}x-2w^{3}y+wxz^{2}-2wyz^{2}+2xz^{3})}{y(z-w)^{3}}\Bigg{)}$
+$\beta^{{\rm alg}}_{\Gamma_{B}}:X(\Gamma_{B})^{{\rm alg}}\longrightarrow\mathbb{P}^{1}_{\mathbb{Q}},\ [x:y:z:w]\mapsto\frac{1}{1728}h\Bigg{(}\frac{z+w}{z-w},\frac{2(w^{3}x-2w^{3}y+wxz^{2}-2wyz^{2}+2xz^{3})}{y(z-w)^{3}}\Bigg{)}$
 is an algebraic Belyi function defined over $\mathbb{Q}$.
+It is also an algebraic model over $\mathbb{Q}$ of $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\simeq\mathbb{P}^{1}(\mathbb{C}),\tau\mapsto\frac{1}{1728}j(\tau)$.
 Remark 2.3.
-The Belyi function $X(G_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\simeq\mathbb{P}^{1}(\mathbb{C})$ in question is a non-Galois covering of $\mathbb{P}^{1}(\mathbb{C})$.
-Thus, it gives rise to a non-regular dessin on $X(G_{B})$.
+The Belyi function $X(\Gamma_{B})\longrightarrow X({\rm SL}_{2}(\mathbb{Z}))\simeq\mathbb{P}^{1}(\mathbb{C})$ in question is a non-Galois covering of $\mathbb{P}^{1}(\mathbb{C})$.
+Thus, it gives rise to a non-regular dessin on $X(\Gamma_{B})$.
 2.3. A decomposition of the Jacobian variety over a number field
-In this section, we study a number field $L$ such that the Jacobian variety ${\rm Jac}(X)$ of $X$ decomposes completely over $L$.
+In this section, we study a number field $L$ such that the Jacobian variety ${\rm Jac}(X(\Gamma_{B})^{{\rm alg}})$ of $X(\Gamma_{B})^{{\rm alg}}$ decomposes completely over $L$.
 For two abelian varieties $A_{1},A_{2}$ over a number field $K$ and a field extension $M/K$, we write $A_{1}\stackrel{{\scriptstyle M}}{{\sim}}A_{2}$ if $A_{1}$ is isogenous over $M$ to $A_{2}$.
@@
 Keep the notation being in Proposition 2.2. Let $L=\mathbb{Q}(\sqrt{5},\sqrt{5+2\sqrt{5}})$. Then, it holds that
-${\rm Jac}(X)\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}A_{f_{{\rm 50.2.b.a}}}\times J_{0}(50)_{\mathbb{Q}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}A_{f_{{\rm 50.2.b.a}}}\times E_{{\rm 50.a3}}\times E_{{\rm 50.b3}}\stackrel{{\scriptstyle L}}{{\sim}}E^{4}$
+${\rm Jac}(X(\Gamma_{B})^{{\rm alg}})\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}A_{f_{{\rm 50.2.b.a}}}\times J_{0}(50)_{\mathbb{Q}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}A_{f_{{\rm 50.2.b.a}}}\times E_{{\rm 50.a3}}\times E_{{\rm 50.b3}}\stackrel{{\scriptstyle L}}{{\sim}}E^{4}$
 where $E_{\ast}$ is the elliptic curve over $\mathbb{Q}$ with conductor 50 whose label $\ast$ corresponds to one in [13] and $E=E_{{\rm 50.a3}}:y^{2}+xy+y=x^{3}-x-2$ is an elliptic curve over $\mathbb{Q}$ with $j$-invariant $-\frac{25}{2}$.
 Proof.
+We refer to
+https://www.lmfdb.org/EllipticCurve/2.2.5.1/100.1/a/3
+linked from
+https://www.lmfdb.org/ModularForm/GL2/Q/holomorphic/50/2/b/a/
+for [13].
 The decomposition over $\mathbb{Q}$ follows from [15, Proposition (2.3)].
-It follows from [13] that $A_{f_{{\rm 50.2.b.a}}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}{\rm Res}_{\mathbb{Q}(\sqrt{5})/\mathbb{Q}}(E_{1})$ where $E_{1}=E_{100.1-{\rm a}3}:y^{2}+(\varepsilon+1)xy+(\varepsilon+1)y=x^{3}+\varepsilon x^{2},\ \varepsilon:=\displaystyle\frac{1+\sqrt{5}}{2}$ with the $j$-invariant $-\frac{25}{2}$.
+It follows from [13] (see the above websites) that $A_{f_{{\rm 50.2.b.a}}}\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}{\rm Res}_{\mathbb{Q}(\sqrt{5})/\mathbb{Q}}(E_{1})$ where $E_{1}=E_{100.1-{\rm a}3}:y^{2}+(\varepsilon+1)xy+(\varepsilon+1)y=x^{3}+\varepsilon x^{2},\ \varepsilon:=\displaystyle\frac{1+\sqrt{5}}{2}$ with the $j$-invariant $-\frac{25}{2}$.
 Here ${\rm Res}_{\mathbb{Q}(\sqrt{5})/\mathbb{Q}}$ stands for the Weil-restriction from $\mathbb{Q}(\sqrt{5})$ to $\mathbb{Q}$.
@@
 4. A relation between two algebraic models
-In this section, we study the field of definition of isomorphisms between $B$, $X$ and the Hulek-Craig curve.
-It is known that Bring’s curve $B$ is isomorphic to the algebraic model $X$ of $X(\Gamma_{B})$ over $\overline{\mathbb{Q}}$.
-We would like to discuss a number field $K$ such that $B$ is isomorphic over $K$ to $X$.
-Obviously, $K$ cannot be $\mathbb{Q}$ since $B(\mathbb{Q})=\emptyset$ while $X(\mathbb{Q})\neq\emptyset$.
+In this section, we study the field of definition of isomorphisms between $B$, $X(\Gamma_{B})^{{\rm alg}}$ and the Hulek-Craig curve.
+It is known that Bring’s curve $B$ is isomorphic to the algebraic model $X(\Gamma_{B})^{{\rm alg}}$ of $X(\Gamma_{B})$ over $\overline{\mathbb{Q}}$.
+We would like to discuss a number field $K$ such that $B$ is isomorphic over $K$ to $X(\Gamma_{B})^{{\rm alg}}$.
+Obviously, $K$ cannot be $\mathbb{Q}$ since $B(\mathbb{Q})$ is empty because of the second quadratic equation in (1.1) while obviously $[1:1:1:0]$ is a $\mathbb{Q}$-rational point on $X(\Gamma_{B})^{{\rm alg}}$.
 For another reason, it is known that ${\rm Jac}(B)\stackrel{{\scriptstyle\mathbb{Q}}}{{\sim}}E^{4}$ (cf.
@@
 Then, $[x:y:z:w]=[w_{1}:w_{2}:w_{3}:w_{4}]A^{-1}$.
-Using this, we have a birational map $\psi$ from the singular model of $X^{{\rm HC}}$ defined by $F(x,y,z)$ to $X$ given by $\psi([x_{1}:y_{1}:z_{1}])=$
+Using this, we have a birational map $\psi$ from the singular model of $X^{{\rm HC}}$ defined by $F(x,y,z)$ to $X(\Gamma_{B})^{{\rm alg}}$ given by $\psi([x_{1}:y_{1}:z_{1}])=$
 (4.1) $\psi([x_{1}:y_{1}:z_{1}])=[x:y:z:w]$
@@
 $\left\{\begin{array}[]{l}x=(z_{1}-y_{1})(x^{2}_{1}+x_{1}y_{1}+y^{2}_{1}+x_{1}z_{1}+z^{2}_{1}),\\ y=(y_{1}-z_{1})(-x^{2}_{1}+x_{1}y_{1}+y^{2}_{1}+x_{1}z_{1}+2y_{1}z_{1}+z^{2}_{1}),\\ z=(y_{1}+z_{1})(-x^{2}_{1}+y_{1}z_{1}),\\ w=x_{1}y^{2}_{1}-y^{3}_{1}+x_{1}z^{2}_{1}-z^{3}_{1}.\end{array}\right.$
-Then, it yields an isomorphism over $\mathbb{Q}$ from $X^{{\rm HC}}$ to $X$. This result is not obvious from the construction explained in [9, p.82-83]. Thus, we have checked the following result:
+Then, it yields an isomorphism over $\mathbb{Q}$ from $X^{{\rm HC}}$ to $X(\Gamma_{B})^{{\rm alg}}$.
+This result is not obvious from the construction explained in [9, p.82-83].
+Thus, we have checked the following result:
 Theorem 4.1.
-Recall that the projective smooth model $X$ over $\mathbb{Q}$ of the modular curve $X(\Gamma_{B})$ from (2.2). It holds that
+Recall that the projective smooth model $X(\Gamma_{B})^{{\rm alg}}$ over $\mathbb{Q}$ of the modular curve $X(\Gamma_{B})$ from (2.2). It holds that
   1. (1)
@@
   2. (2)
-    Hulek-Craig’s curve $X^{{\rm HC}}$ is isomorphic over $\mathbb{Q}$ to $X$ whose isomorphism is given by (4.1).
+    Hulek-Craig’s curve $X^{{\rm HC}}$ is isomorphic over $\mathbb{Q}$ to $X(\Gamma_{B})^{{\rm alg}}$ whose isomorphism is given by (4.1).
 In particular, Hulek-Craig’s curve is isomorphic over $\mathbb{Q}$ to a canonical model of $X(\Gamma_{B})$ (see [11] for canonical models of modular curves).
-Thus, $B$ inherits two algebraic Belyi functions over $\mathbb{Q}(\zeta_{5})$ such that one is non-regular (not a Galois covering) while the other is regular (a Galois covering) (see [7] for the definition of regular Belyi functions).
+Thus, $B$ inherits two algebraic Belyi functions such that one is defined over $\mathbb{Q}(\zeta_{5})$ and non-regular (not a Galois covering) while the other is defined over $\mathbb{Q}$ and regular (a Galois covering) (see [7] for the definition of regular Belyi functions).